the price of a winter jacket was decreased to a sale of $56.00. this was a 30% reduction.what was the former price?
Answer by richwmiller(17219)

.70*x=56
7x/10=56
7x=560
x=80
80 dollars was the original price
check
.7*80=56
7/10*80=56
7*8=56
ok
3/10*80=24
80-24=56
ok